OpenStudy (anonymous):

differentiation question: draw the curve y=ln(x+1) for x values between 0 & 4 included. by drawing an appropriate straight line, obtain an approximate value for the positive square root of the equation x-2ln(x+1)=0.

OpenStudy (anonymous):

i can't say i understand what this question wants. especially the part of about the "positive square root of \[x-2\ln(x+1)\] which frankly makes no sense

OpenStudy (anonymous):

however, if you are trying to solve \[x-\ln(x+1)=0\] for x, then \[x=0\] works by inspection

OpenStudy (anonymous):

makes sense, but what about the second part?

OpenStudy (anonymous):

that is the second part

OpenStudy (anonymous):

isnt that just x-2y=0

OpenStudy (anonymous):

here is the picture of \[y=\ln(x+1)\] on the interval [0,4] http://www.wolframalpha.com/input/?i=ln%28x%2B1%29+x%3D0+to+4

OpenStudy (anonymous):

I used wolfram too :P but what i dont get is the positive square root of the equation. how do i find that?

OpenStudy (amistre64):

primary square root maybe?

OpenStudy (anonymous):

what is this "square root" business?! it is not a square

OpenStudy (amistre64):

\[\pm sqrt{...}, \color{#dd0011}{\text{use the +}}\]

OpenStudy (amistre64):

dunno the business part lol

OpenStudy (anonymous):

\[x-2\ln(x+1)=0\] has two solutions, but in no sense are they "square roots" of anything

OpenStudy (amistre64):

\[ \sqrt{x-2ln(x+1)};\ x-2ln(x+1)\ge0\]

OpenStudy (anonymous):

and how this relates to the first question is beyond me. i thought it was going to be a visualization of newton's method, but it is not. the first expression is not the same as the second one

OpenStudy (anonymous):

maybe it is supposed to read "obtain an approximate value for the positive root of the equation x-2ln(x+1)=0. "

OpenStudy (amistre64):

approximate: ln(x+1) with an appropriate line ... yada yada

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

yeah but that doesn't help solve \[x-2\ln(x+1)=0\] as far as i can see

OpenStudy (amistre64):

the equation of the tangent line to ln(x+1) at x gives as a close approx of ln(x+1)

OpenStudy (amistre64):

maybe we need to use the approximation of ln(x+1) in its place

OpenStudy (amistre64):

or maybe its just the average rate of change from 0 to 4 that we use

OpenStudy (amistre64):

http://openstudy.com/?F225231636612OYKJOD=_#/updates/4eec41e9e4b0367162f56dc3 might be more info here

OpenStudy (anonymous):

no, those are separate questions

OpenStudy (amistre64):

y = .4024 x is the line drawn from 0 to 4 http://www.wolframalpha.com/input/?i=y+%3D+%28ln%285%29-ln%281%29%29x%2F%284%29and+ln%28x%2B1%29+x%3D0+to+4

OpenStudy (amistre64):

if we use that in place of ln(x+1) we can get an approx of the equation instead of having to use ln(x+1)

OpenStudy (amistre64):

you can get a better and better approx by using like the first 5 terms of its taylor series expansion

OpenStudy (anonymous):

oh okay thanks guys!

OpenStudy (amistre64):

thats the best I can make sense of it, good luck with it :)